OpenStudy (anonymous):

optimization problem: a physical fitness room consists of a rectangular region with a semicircle on each end. if the perimeter of the room is to be a 200 meter running track, find the dimensions that will make the are of the rectangular region as large as possible

OpenStudy (anonymous):

Do you know how to find the largest area possible?

OpenStudy (anonymous):

Draw a picture

OpenStudy (anonymous):



OpenStudy (anonymous):

lol

OpenStudy (anonymous):

That's fine lol

OpenStudy (anonymous):

c=2piR and 200=4r+2x (x being the base of the rectangle) this is how i set it up

OpenStudy (anonymous):

i think you have the perimeter wrong

OpenStudy (anonymous):

how?

OpenStudy (anonymous):

You're not accounting for the circumference of the circle. Your equation only includes the area of the rectangle.

OpenStudy (anonymous):

You would need it to be \(2x + c = 200\)

OpenStudy (anonymous):

c is the circumference right

OpenStudy (anonymous):

Yes, and that's right (\(2 \pi r\)).

OpenStudy (anonymous):

okay and then differentiate that equation correct?

OpenStudy (anonymous):

You would need to differentiate the area, not the perimeter.

OpenStudy (anonymous):

So, can you first determine a formula for the area of your diagram?

OpenStudy (anonymous):

pir^2+2rx

OpenStudy (anonymous):

Yeah, that looks good.

OpenStudy (anonymous):

should i substitute r or x

OpenStudy (anonymous):

Whichever one you want

OpenStudy (anonymous):

100/3pi = r seem right?